# Tracking Brownian fluid particles in large eddy simulations

### v2

@@ 4.2 Experiments for laminar flows
Now we consider the bounded case. In this part, $\nu=0.3$, $Re=1000$, length scale $L_{h}=3\pi,L_{v}=0.2\pi$, so that $U_{0}=\frac{\nu}{L_{h}}Re=31.83$. The mesh size $s_{h}=\frac{3\pi}{50},s_{v}=\frac{0.2\pi}{50}$. and the time step $\delta=0.001$. The numerical experiment is demonstrated at times $t=0.03,t=0.06,t=0.09$. We set the initial velocity to be of the form $U(x,0)=(U_{0},0)$, and set force $F=(10e^{(-\frac{(x_{1}^{2}+x_{2}^{2})}{2s_{h}s_{v}})},-9.81)$. The velocity field and the vorticity field are shown in Figure 1.
@@
Figure 2: Velocity fields (left) and vorticity (right) of wall-bounded laminar flows on $\mathbb{R}_{+}^{2}$
@@
As shown in Figure 2, the fluid system initially remains in a relatively stable state. Over time, a small number of vortices begin to emerge at the boundary and gradually propagate into the interior of the domain. This phenomenon is primarily driven by shear forces, specifically the combined effects of boundary pressure and nonlinear forces. Notably, compared to the subsequent turbulent flow experiments (Figure 12), the fluid dynamics in the laminar flow experiments exhibit significantly lower intensity, with fewer vortices generated. This observation is consistent with established theoretical principles.
4.2.3 3-dimensional unbounded laminar flows
Similar to 2-dimensional numerical experiments, we set $\nu=0.3$, $Re=1000$, length scale $L=3\pi$, so that $U_{0}=\frac{\nu}{L}Re=31.83$. The mesh size $s=\frac{3\pi}{50}\thicksim L\sqrt{\frac{1}{Re}}$. and the time step $\delta=0.001$. The numerical experiment is demonstrated at times $t=0.03,t=0.06,t=0.09$. We set the initial velocity to be of the form $U(x,0)=(U_{0}\sin(2x_{2}),U_{0}\cos(2x_{1}),0)$, and set force $F=(10,10,-9.81)$. The velocity field and the vorticity field are shown in Figure 3.
Figure 3: Velocity fields (left) and vorticity (right) of unbounded laminar flows on $\mathbb{R}^{3}$
Similar to the unbounded two-dimensional experiment, the initial rotating velocity field undergoes a transition from its initial formation to gradual dissipation. As shown in Figure 3, the initial conditions generate a vortex-like flow pattern, while the vorticity exhibits increasingly chaotic behavior. To enhance the visualization of the flow dynamics, sectional views of the velocity fields are presented at the planes $x=C$, $y=C$, and $z=C$ in Figures 4, 5, and 6, respectively, with $C=0.1\pi$ chosen for illustration.
Figure 4: Section Velocity fields of incompressible viscous flows on plane on $x=C$
Figure 5: Section Velocity fields of incompressible viscous flows on plane on $y=C$
Figure 6: Section Velocity fields of incompressible viscous flows on plane on $z=C$
The sectional velocity fields are influenced by the initial conditions specified in the setup. Based on the initial velocity field configuration, the flow on $z=C$ plane exhibits more complex behavior, as illustrated in Figure 6. From the sectional views, it is observed that as time progresses, the fluid predominantly moves in the direction of the applied force, consistent with the expected physical behavior.
4.2.4 3-dimensional wall-bounded laminar flows
In this experiment, $\nu=0.3$, $Re=1000$, $L_{h}=3\pi,L_{v}=0.2\pi$, so that $U_{0}=\frac{\nu}{L}Re=31.83$. The horizontal mesh size $s_{h}=\frac{3\pi}{50}$, vertical mesh size $s_{v}=\frac{0.2\pi}{50}$ , and the time step $\delta=0.001$. The numerical experiment is demonstrated at times $t=0.03,t=0.06,t=0.09$. We set the initial velocity to be of the form $U(x,0)=(U_{0},U_{0},0)$, and set force $F=(100e^{(-\frac{(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})}{2s_{h}^{2}s_{v}})},100e^{(-\frac{(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})}{2s_{h}^{2}s_{v}})},-9.81)$. The velocity field and the vorticity field are shown in Figure 7.
Figure 7: Velocity fields (left) and vorticity (right) of wall-bounded laminar flows on $\mathbb{R}_{+}^{3}$
In contrast to the unbounded case, the presence of gravity and boundary shear forces induces fluctuations in the flow near the boundary, leading to the formation of vorticity. As a result, the absolute value of the boundary vorticity increases, as shown in Figure 7. We find that the formation of the vortex is not so complete, which is because the changes in laminar flows are not so intense.
Figure 8: Section Velocity fields of incompressible viscous flows on plane on $x=C$
Figure 9: Section Velocity fields of incompressible viscous flows on plane on $y=C$
Figure 10: Section Velocity fields of incompressible viscous flows on plane on $z=C$
The three sectional figures illustrate that on the tangent plane adjacent to the boundary, fluctuations and variations occur to some extent. However, these changes are not significant due to the nature of the laminar flow. Moreover, the fluid flow on the $z=C$ plane largely maintains its initial state.
4.3 Experiments for turbulent flows
4.3.1 2-dimensional unbounded turbulent flows
In this subsection, we present a numerical experiment involving a 2-dimensional turbulent flow in $\mathbb{R}^{2}$. For this experiment, we set the kinematic viscosity $\nu=0.3$ and the Reynolds number $Re=5500$. The horizontal and vertical length scales ($L_{h}$ and $L_{v}$) are both set to $L_{h}=L_{v}=3\pi$, From the relationship between velocity and the Reynolds number, we calculate the initial mean stream velocity as $U_{0}=\frac{\nu}{L}Re=135.15$. The computational grid is defined with a horizontal mesh size $s_{h}=\frac{3\pi}{50}\thicksim L\sqrt{\frac{1}{Re}}$ and a vertical mesh size $s_{v}=\frac{3\pi}{50}$. The time step is set to $\delta=0.0001$. The numerical results are presented at three time instances: $t=0.01,t=0.02,t=0.03$. The initial velocity field is specified as $U(x,0)=(U_{0}sin(2x_{2}),U_{0}cos(2x_{1}))$ and the external force is given by $F=(10,-9.81)$. The velocity fields are shown in Figure 11 (left), and the corresponding vorticity is displayed in Figure 11 (right).
Figure 11: Velocity fields (left) and vorticity (right) of unbounded turbulent flows on $\mathbb{R}^{2}$
Since no external boundary conditions or nonlinear forces were applied in this experiment, an initial rotating vector field is required to observe the system’s evolution. This rationale underpins the chosen form for the initial velocity field. As shown in Figure 11, the velocity field evolves from an initial rotating pattern into a more irregular state, eventually transitioning into a chaotic regime. Correspondingly, the vorticity exhibits fluctuations, consistent with the expected behavior based on our experimental setup and hypothesis.
4.3.2 2-dimensional wall-bounded turbulent flows
@@ 4.3 Experiments for turbulent flows
In this subsection, we present a numerical experiment simulating a two-dimensional turbulent flow on $\mathbb{R}_{+}^{2}$, where the flow passes over a solid wall at a high Reynolds number. To better capture the boundary effects, we use a smaller observation domain and finer spatial grid resolution in the vertical direction. The parameters for the experiment are as follows: $\nu=0.3$, $Re=5500$, horizontal length scale $L_{h}=3\pi$, vertical length scale $L_{v}=0.2\pi$ and the initial main stream velocity $U_{0}=135.15$. The horizontal mesh size $s_{h}=\frac{3\pi}{50}$, vertical mesh size $s_{v}=\frac{0.2\pi}{50}$ and the time step $\delta=0.0001$. The numerical results are demonstrated at three time instances: $t=0.01,t=0.02,t=0.03$. The initial velocity field is defined as $U(x,0)=(U_{0},0)$ while the external force is given by $F=(100e^{(-\frac{(x_{1}^{2}+x_{2}^{2})}{2s_{h}s_{v}})},-9.81)$. The velocity and vorticity are shown in Figure 12.
@@
Figure 12: Velocity fields (left) and vorticity (right) of wall-bounded turbulent flows on $\mathbb{R}_{+}^{2}$
@@
As shown in Figure 12, the fluid system initially remains in a relatively stable state. Over time, vortices begin to form at the boundary and gradually propagate into the interior of the domain. This phenomenon is primarily attributed to shear forces, specifically the combined influence of boundary pressure and nonlinear forces. In contrast to the boundary-free case, the presence of a solid boundary inherently facilitates vortex generation at the wall. Additionally, the vorticity near the boundary exhibits significantly higher values, which aligns with physical observations in real-world fluid flows.
4.3.3 3-dimensional unbounded turbulent flows
Numerical experiment simulating a three-dimensional turbulent flow at a high Reynolds number is shown in this subsection. In this experiment, the kinematic viscosity is set to $\nu=0.3$ and the Reynolds number is $Re=5500$. Similar to the two-dimensional simulations, the characteristic length scales in both directions are $L_{h}=L_{v}=3\pi$, which gives an initial mean stream velocity of $U_{0}=135.15$. The mesh size $s=\frac{3\pi}{50}$, and the time step $\delta=0.0001$. Numerical results are presented at the following time instances: $t=0.01,t=0.02,t=0.03$. The initial velocity field is defined as $U(x,0)=(U_{0}\sin(2x_{2}),U_{0}\cos(2x_{1}),0)$, and the external force is given by $F=(100,100,-9.81)$. Figure 17 shows the results of the experiment.
Figure 13: Velocity fields (left) and vorticity (right) of unbounded turbulent flows on $\mathbb{R}^{3}$
Similar to the unbounded two-dimensional experiment, in the three-dimensional case, we must either set the initial velocity field as a rotating vector field or add the nonlinear force to observe fluid dynamics. In this experiment, we choose the former approach. As shown in Figure 13, the initial conditions result in a vortex-like flow pattern, while the vorticity exhibits more chaotic behavior. Additionally, as time progresses, the fluid generally flows in the direction of the applied force, which aligns with expected physical behavior. To visualize the flow more clearly, we provide sectional views of the velocity fields at planes $x=C$, plane $y=C$, and the plane $z=C$ in Figure 14, Figure 15, Figure 16 respectively (we choose $C=0.1\pi$).
Figure 14: Section Velocity fields of incompressible viscous flows on plane on $x=C$
Figure 15: Section Velocity fields of incompressible viscous flows on plane on $y=C$
Figure 16: Section Velocity fields of incompressible viscous flows on plane on $z=C$
As shown in Figures 14, 15, and 16, the fluid system gradually becomes chaotic and disordered on the $x=C$ and $y=C$ planes, as only linear forces and trivial initial conditions are applied in these two directions. In contrast, on the $z=C$ plane, the vortex phenomenon is more pronounced due to the initial rotational conditions specified in this plane.
4.3.4 3-dimensional wall-bounded turbulent flows
In this subsection, we present a numerical experiment simulating a three-dimensional turbulent flow in $\mathbb{R}_{+}^{3}$, where the flow passes over a solid wall at a high Reynolds number. To better capture the boundary effects, we use a finer spatial grid resolution in the vertical direction. The experiment is conducted with the following parameters: the kinematic viscosity $\nu=0.3$, $Re=5500$, horizontal length scale $L_{h}=3\pi$, vertical length scale $L_{v}=0.2\pi$, initial main stream velocity $U_{0}=135.15$. The horizontal mesh size $s_{h}=\frac{3\pi}{50}$, vertical mesh size $s_{v}=\frac{0.2\pi}{50}$and the time step $\delta=0.0001$. The numerical experiment results are demonstrated at times $t=0.01,t=0.02,t=0.03$. The initial velocity field is given by $U(x,0)=(U_{0},U_{0},0)$, set force is given by $F=(100e^{(-\frac{(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})}{2s_{h}^{2}s_{v}})},100e^{(-\frac{(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})}{2s_{h}^{2}s_{v}})},-9.81)$. The velocity and vorticity are shown in Figure 17.
Figure 17: Velocity fields (left) and vorticity (right) of wall-bounded turbulent flows on $\mathbb{R}_{+}^{3}$
Figure 17 shows that, in the 3D turbulent flow, the velocity field evolves over time, with vortices gradually forming at the boundary $z=0$. The vorticity field exhibits corresponding changes, which aligns with both theoretical predictions and real-world observations. To better visualize the flow, we provide cross-sectional views of the velocity field at the planes $x=C$, $y=C$, and $z=C$ in Figures 18, 19, and 20, respectively, where $C=0.1\pi$.
@@
In Figures 18 and 19, the application of external forces leads to the formation of eddies at the boundary, which then propagate throughout the domain as the fluid flows across the wall surface. In contrast, on the $z=C$ plane (Figure 20), where no initial nonlinear forces are applied in this direction, only minor changes in the flow are observed, consistent with the initial conditions we set. These experimental results are in agreement with both theoretical predictions and empirical observations, thereby validating our model and numerical scheme.
In general, compared to laminar flow, turbulent flow exhibits a non-uniform velocity distribution with more pronounced gradient variations. Furthermore, turbulent flow is characterized by significant energy dissipation and inherent instability. It generates a greater number of vortices near the wall boundary, which tend to be more fully developed in structure. In contrast, laminar flow produces fewer vortices, which may not fully develop or maintain a complete structure. The real-world phenomena described above are effectively replicated in our experiments, validating the effectiveness and reliability of our methodology.
5 Appendix: Fluid flows represented by Brownian fluid particles

### v4

@@ 4.2.1 Experiment 1: 2-dimensional advection through a wall (laminar flows)
Now we consider the bounded case. In this part, $\nu=0.3$, $Re=1000$, length scale $L_{h}=3\pi,L_{v}=0.2\pi$, so that $U_{0}=\frac{\nu}{L_{h}}Re=31.83$. The mesh size $s_{h}=\frac{3\pi}{50},s_{v}=\frac{0.2\pi}{50}$. and the time step $\delta=0.001$. The numerical experiment is demonstrated at times $t=0.03,t=0.06,t=0.09$. We set the initial velocity to be of the form $U(x,0)=(U_{0},0)$, and set force $F=(10e^{(-\frac{(x_{1}^{2}+x_{2}^{2})}{2s_{h}s_{v}})},-9.81)$. The velocity field and the vorticity field are shown in Figure 1 and 2.
@@
Figure 1: Velocity fields of wall-bounded laminar flows on $\mathbb{R}_{+}^{2}$
Figure 2: Vorticity of wall-bounded laminar flows on $\mathbb{R}_{+}^{2}$
@@
As shown in Figure 1, the fluid system initially remains in a relatively stable state. Over time, a small number of vortices begin to emerge at the boundary and gradually propagate into the interior of the domain. This phenomenon is primarily driven by shear forces, specifically the combined effects of boundary pressure and nonlinear forces. Notably, compared to the subsequent turbulent flow experiments (Figure 3), the fluid dynamics in the laminar flow experiments exhibit significantly lower intensity, with fewer vortices generated.
In laminar or low Reynolds number case, where viscous forces prevail over fluid inertia, the vorticity dynamics are characterized by diffusion and wall confinement. The no-slip condition at the wall ($y=0$) acts as a source of strong vorticity. However, its transport into the flow interior occurs primarily through slow molecular diffusion, leading to rapid attenuation of strength with distance from the wall. The observed incomplete vortices are, in essence, starting or separation vortices formed by the rolling-up of this diffused vorticity layer under the action of external shear. The fundamental reason for their distorted morphology is twofold: the concentration of vorticity is insufficient, and the rotational energy of the nascent vortex is too weak to overcome the powerful viscous dissipation and the anchoring effect of the wall. Consequently, these structures remain tethered to the boundary, unable to lift off and coalesce into a closed, symmetric vortex core. Their vorticity is primarily sustained by the local wall shear rate rather than by large-scale vortex entrainment. This observation is consistent with established theoretical principles.
4.2.2 Experiment 2: 2-dimensional advection through a wall (turbulent flows)
@@
In this subsection, we present a numerical experiment simulating a two-dimensional turbulent flow on $\mathbb{R}_{+}^{2}$, where the flow passes over a solid wall at a high Reynolds number. To better capture the boundary effects, we use a smaller observation domain and finer spatial grid resolution in the vertical direction. The parameters for the experiment are as follows: $\nu=0.3$, $Re=5500$, horizontal length scale $L_{h}=3\pi$, vertical length scale $L_{v}=0.2\pi$ and the initial main stream velocity $U_{0}=135.15$. The horizontal mesh size $s_{h}=\frac{3\pi}{50}$, vertical mesh size $s_{v}=\frac{0.2\pi}{50}$ and the time step $\delta=0.0001$. The numerical results are demonstrated at three time instances: $t=0.01,t=0.02,t=0.03$. The initial velocity field is defined as $U(x,0)=(U_{0},0)$ while the external force is given by $F=(100e^{(-\frac{(x_{1}^{2}+x_{2}^{2})}{2s_{h}s_{v}})},-9.81)$. The velocity and vorticity are shown in Figure 3 and 4.
@@
Figure 3: Velocity fields of wall-bounded turbulent flows on $\mathbb{R}_{+}^{2}$
Figure 4: Vorticity of wall-bounded turbulent flows on $\mathbb{R}_{+}^{2}$
@@
As shown in Figure 3 and 4, the fluid system initially remains in a relatively stable state. Over time, vortices begin to form at the boundary and gradually propagate into the interior of the domain. This phenomenon is primarily attributed to shear forces, specifically the combined influence of boundary pressure and nonlinear forces. Additionally, the vorticity near the boundary exhibits significantly higher values, which is consistent with physical observations in real-world fluid flows.
In the turbulent regime, inertial effects dominate. The wall remains the primary source of vorticity, but turbulence greatly enhances its production and transport through stretching and reorientation mechanisms. Small-scale turbulent fluctuations tear and lift the wall-generated vorticity layers, rolling them into concentrated vortex tubes or packets. Compared with laminar flows (Experiment 1), the complete vortices observed in turbulent flow owe their integrity to the sufficient kinetic energy contained within their rotational cores, which enables them to lift away from the wall and extend into the outer flow region. These vortices form relatively closed streamline loops with clearly identifiable low-pressure cores, indicating a well-developed and coherent vortex structure. The larger magnitude of vorticity near the turbulent boundary does not arise from an infinitely increasing rate of wall generation, but rather from the intensified transport, accumulation, and amplification of vorticity driven by turbulent motions.
4.2.3 Experiment 3: 3-dimensional initially stationary fluid
In this subsection, we set an interesting 3-dimensional experiment. We assume that the flow is stationary initially (which is actually a laminar flow), then a tensile force in the same direction was applied to the fluid, observing the changes in the fluid system. We set $\nu=0.3$, length scale $L=3\pi$, the mesh size $s=\frac{3\pi}{50}\thicksim L\sqrt{\frac{1}{Re}}$ and the time step $\delta=0.001$. The numerical experiment is demonstrated at times $t=0.1,t=0.2,t=0.3$. We set the initial velocity to be of the form $U(x,0)=(0,0,0)$, and set force $F=(100e^{(-\frac{(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})}{2s_{h}^{2}s_{v}})},100e^{(-\frac{(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})}{2s_{h}^{2}s_{v}})},-9.81)$. The velocity field and the vorticity field are shown in Figure 5 and 6.
Figure 5: Velocity fields wall-bounded flows on $\mathbb{R}^{3}_{+}$
Figure 6: Vorticity of wall-bounded flows on $\mathbb{R}^{3}_{+}$
Figure 5 shows that the initially stationary fluid system flows along the direction of the pulling force, and as time progresses, vortices are generated at the boundaries. When a tensile force is suddenly applied to an initially stationary (laminar) fluid, the system experiences a strong streamwise acceleration that establishes a velocity gradient between the bulk and the no-slip boundaries. This gradient immediately generates vorticity at the walls, where the velocity remains zero. The wall-bounded vorticity diffuses into the fluid and is simultaneously advected by the emerging mean flow. As time evolves, the non-uniform acceleration induces spatial variations in the strain rate, producing regions of local vorticity amplification. The interaction between the streamwise stretching and the wall-generated shear promotes vortex formation near the boundaries, as observed in figure 5. These vortices form from the partial roll-up of wall-generated vorticity layers under the applied extensional flow. Due to the relatively low flow velocity, the rotational energy is limited, resulting in irregular and asymmetric vortex shapes.
In order to better depict the shape of the flow, we make two-dimensional velocity field slices to visually represent the behavior of the fluid. The sectional views are presented at the planes $x=C$, $y=C$, and $z=C$ in Figures 7, 8, and 9, respectively, with $C=0.1\pi$ chosen for illustration.
@@ 4.2.3 Experiment 3: 3-dimensional initially stationary fluid
With the tensile force applied only in the $xy$-plane, the fluid accelerates primarily along $x$ and $y$, generating velocity gradients near the wall at $z=0$. These vertical gradients produce vorticity oriented along the $z$-direction, while horizontal slices in the $xy$-plane show little to no vortex activity. As the flow develops, the wall-generated vorticity is stretched and deformed along the $xy$-plane, forming irregular, elongated vortices confined near the boundaries, reflecting the anisotropic forcing and limited vertical transport, which is consistent with the expected physical behavior.
4.2.4 Experiment 4: 3-dimensional advection through a wall (turbulent flows)
In order to further demonstrate the effectiveness of our method, we complete the experiment of initial advection turbulent advection passing through a wall. In this experiment, $\nu=0.3$, $Re=5500$, $L_{h}=3\pi,L_{v}=0.2\pi$, so that $U_{0}=\frac{\nu}{L}Re=135$. The horizontal mesh size $s_{h}=\frac{3\pi}{50}$, vertical mesh size $s_{v}=\frac{0.2\pi}{50}$ , and the time step $\delta=0.001$. The numerical experiment is demonstrated at times $t=0.1,t=0.2,t=0.3$, and we also show the structure of flows in $t=0.01$ to see the behavior of the fluid in the initial stage. We set the initial velocity to be of the form $U(x,0)=(U_{0},U_{0},0)$, and set force $F=(100e^{(-\frac{(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})}{2s_{h}^{2}s_{v}})},100e^{(-\frac{(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})}{2s_{h}^{2}s_{v}})},-9.81)$. The velocity field and the vorticity field are shown in Figure 10 and 11.
Figure 10: Velocity fields of wall-bounded turbulent flows on $\mathbb{R}_{+}^{3}$
Figure 11: Vorticity fields of wall-bounded turbulent flows on $\mathbb{R}_{+}^{3}$
In this experiment, we specifically present fluid flow visualizations over a short time scale. At early times ($t=0.01$), the flow remains close to its initial advective state, with negligible vorticity away from the walls. As the flow evolves, shear near the boundary walls generates strong wall-bounded vorticity due to the no-slip condition, which is initially confined close to the wall. This vorticity is amplified and stretched by the developing flow, producing vortices that grow in size and complexity, as illustrated in Figures 10 and 11.
Compared with laminar flows, inertial effects in the turbulent regime dominate, leading to highly three-dimensional and unsteady dynamics. The enhanced advection and vortex stretching in turbulence increase the local vorticity magnitude near the walls, making boundary-layer vortices more intense than in laminar cases. Small-scale turbulent fluctuations continuously distort and redistribute the wall-generated vorticity, giving rise to irregular, asymmetric vortex structures that penetrate into the outer flow. In effect, the observed vorticity field reflects the combined influence of wall generation, inertial amplification, and three-dimensional turbulent transport. To enhance the clarity of the flow structures, two-dimensional sectional views are shown.
In plane $x=C$ (Figure 12) and $y=C$ (Figure 13), the overall flow direction of the fluid remains largely aligned with the initial advective direction and the direction of the applied force. As time progresses, fully developed vortices gradually emerge from the boundaries. The explanation of the flow mechanism of the section is similar to that of Experiment 3.
4.3 Analysis and comparison
4.3.1 Divergence-free analysis
For 2D flows, the value of divergence is easy for visualization. But for 3D case, the divergence is not easily visualized through 3D images, the absolute value of the divergence at each grid point is calculated and then the spatial average is taken. We print the time series figures of the overall average divergence of the fluid system. Following the conventions in the existing literature, a regular method to verify the divergence-free analysis is to calculate whether $\frac{|\nabla\cdot u|}{|\nabla u|}\ll 1$. To achieve this, we define
[…]
to calculate the mean divergence value of each timestep, the results are shown in Figure 15 and 16.
(a) Experiment 1
(b) Experiment 2
Figure 15: Incompressible test for Experiment 1 (left) and Experiment 2 (right)
From Figure 15, it can be observed that $\frac{|\nabla\cdot u|}{|\nabla u|}$ is significantly less than 1 in the two-dimensional case. A comparison between Experiment 1 and Experiment 2 reveals that the enforcement of the incompressibility condition is less effective in Experiment 2, owing to the presence of turbulence at high Reynolds numbers. The results indicate that the Random LES method generally preserves the incompressibility assumption.
(a) Experiment 3
(b) Experiment 4
Figure 16: Incompressible test for Experiment 3 (left) and Experiment 4 (right)
For more complex three-dimensional cases, the incompressibility assumption remains valid. Moreover, fluid systems with higher velocities tend to exhibit larger values of $\text{test}(t)$ than systems with lower velocities, consistent with the trends observed in the two-dimensional cases.
4.3.2 Comparison with other methods
The method proposed in this study is relatively novel and offers considerable scope for further optimization. Consequently, in this subsection, we compare and analyze our method with other available methods to identify the advantages of our approach. However, the primary objective of this work is to introduce the new approach and establish its validity, rather than to compare it against fully optimized and well-established methods, so we compare our method with early versions of other methods. To provide a preliminary benchmark, two classical numerical techniques are employed: the finite difference method (FDM) and the finite volume method (FVM), both of which are widely adopted and critically important in computational fluid dynamics. The FDM approximates the derivative terms in differential equations using finite differences, constructing discrete equations from function values at grid nodes. In contrast, the FVM divides the computational domain into a set of control volumes and derives discrete equations by integrating the governing differential equations over each control volume. What’s more, we also compare Random LES method with mild solution method mentioned in Section 1, which is widely studied in the theory of stochastic PDE. Using Experiment 4 as an illustrative example, we present an analysis comparing the performance of these methods. All simulations are conducted on Intel(R) i7-12700H CPU (2.30 GHz) .
For FDM method, we obey the assumption that $\Delta t\leq\frac{\Delta x}{|U_{m}ax|}$, which means that for small mesh size, we should select a small timestep $\Delta t$ to enable the stability of numerical method. For the FVM method, we use the package in python to finish the simulation. Then we add the simulation utilizing the mild solution mentioned in Section 1, the formula is:
$u^{i}(x,t)-u^{i}_{0}(x)=$
$\int_{0}^{t}\int_{D}\left(u^{i}(y,s)u(y,s)\cdot\nabla_{y}\ln{h}(x,t-s,y)-{\partial\over\partial y_{i}}p(y,s)-F^{i}(y,s)\right)h(x,t-s,y)\textrm{d}y\textrm{d}s,$
for $i=1,2,3$. The velocity $u(x,t)$ is updated by the variables in time t1t-1, the calculation of $\nabla p$ is calculated by the same method (kernel method) in Random LES and the gradient of $u$ needs to be calculated by FDM method since we can not gain the explicit expression of it. We detect the stability of all the numerical methods by observing when the numerical solution blows up. The results are shown in Table 1.
Table 1: Numerical instability test
Random LES method FDM method FVM method Mild solution
$\Delta x=\Delta y=\frac{3\pi}{50}$,
$\Delta z=\frac{0.2\pi}{50}$ , $\Delta t=0.01$
Stable Explode at step 10 Stable Explode at step 22
$\Delta x=\Delta y=\frac{3\pi}{50}$,
$\Delta z=\frac{0.2\pi}{50}$ , $\Delta t=0.001$
Stable Explode at step 21 Stable Explode at step 45
$\Delta x=\Delta y=\frac{3\pi}{100}$,
$\Delta z=\frac{0.2\pi}{100}$ , $\Delta t=0.001$
Stable Explode at step 41 Stable Explode at step 75
It is evident that, without the incorporation of complex boundary constraints, both the FDM and mild solution methods are prone to numerical instability and potential blow-up. The latter tends to diverge more slowly, as it involves only first-order derivative approximations and avoids the computation of second-order derivatives. In contrast, the Random LES method and FVM demonstrate greater stability. While numerous researchers continue to explore various optimizations for the FDM, our analysis focuses on comparisons with its baseline versions. Theoretically, our Monte Carlo-based approach is less susceptible to numerical instability. Subsequently, we evaluated the average runtime per epoch, with the results presented in Table 2.
Table 2: Average calculation time of each epoch
Random LES method FDM method FVM method Mild solution
$\Delta x=\Delta y=\frac{3\pi}{50}$,
$\Delta z=\frac{0.2\pi}{50}$, $\Delta t=0.01$
4.69 seconds 0.449 seconds 6.5 seconds 26.5 seconds
$\Delta x=\Delta y=\frac{3\pi}{50}$,
$\Delta z=\frac{0.2\pi}{50}$ , $\Delta t=0.001$
7.71 seconds 1.143 seconds 13 seconds 73.4 seconds
$\Delta x=\Delta y=\frac{3\pi}{100}$,
$\Delta z=\frac{0.2\pi}{100}$ , $\Delta t=0.001$
16.5 seconds 11.54 seconds 78.7 seconds 165.4 seconds
In terms of computational time, the FDM, benefiting from its integration into existing software packages, generally achieves faster computations compared to our approach. Nonetheless, we argue that the computational cost of our method remains well within acceptable limits. More importantly, as grid resolution increases, FDM’s computational time escalates significantly and is accompanied by an increased risk of numerical instability. In contrast, our method exhibits stable computational scalability without these limitations. The mild solution method requires the longest computational time among the compared methods, with runtime increasing substantially as mesh resolution becomes finer. Because the kernel term and the other terms are time-reverse, which indicates that we need to recalculate integral from 0 to $t$ in every iterative step in mild solution method, which is very expensive in calculation. Finally, we present a theoretical analysis of the computational complexity of the proposed numerical scheme.
Table 3: Analysis of complexity
Updating variable Random LES method FDM method FVM method Mild solution
Brownian particle $Y_{t}$ Using value at t1t-1 \ \ \
Velocity $U_{t}$ Using 1-order derivative Using 2-order derivative Using 2-order derivative Using 1-order derivative
Pressure $P$ \ Using 2-order derivative Using 2-order derivative \
Gradient of $P$ Using 1-order derivative \ \ Using 1-order derivative
We observe that the numerical scheme of the Random LES method requires only the computation of first-order derivatives, while both the FDM and FVM involve second-order derivatives. Theoretically, the Random LES approach offers distinct advantages: calculating second-order derivatives not only leads to greater error accumulation but also increases the risk of numerical instability and entails higher computational cost, particularly in high-dimensional settings. Although the mild solution method relies solely on first-order derivatives, the term $\nabla u$ cannot be explicitly expressed. Therefore, we still need to employ additional numerical techniques to approximate it.
According to the results in Table 1, 2, 3 and the corresponding analysis, compared to alternative approaches, the Random LES method generally demonstrates superior stability, acceptable computational efficiency, and relatively lower theoretical complexity. A key advantage of our approach lies in its ability to reveal diverse potential structures of the fluid system through individual simulations. Since the exact morphology of the fluid cannot be perfectly determined through computation, conventional deterministic methods that produce fixed solutions may fail to capture significant fluid behaviors emerging from our probabilistic framework. Our Random LES method, which is based on explicit variable representation and probabilistic format, significantly improves the accuracy and stability of numerical experiments without substantially increasing computational time.
Appendix A Fluid flows represented by Brownian fluid particles